OpenStudy (anonymous):

The lengths of the legs of a right triangle are 6 inches and 8 inches. What is the length of the hypotenuse?

OpenStudy (anonymous):

6^2 + 8^2 = hypotnuse^2 36+64 =100 sqrt of 100 = 10

OpenStudy (anonymous):

Hypotenuse is the greatest side of all in a right triangle.. So, you have given two sides, use pythagoras theorem here: \[\large Hypotenuse^2 = a^2 + b^2\] Here a and b are the other sides: \[H^2 = 6^2 + 8^2 \implies H^2 = 36 + 64 \implies H^2 = 100\] Take square root now: \[\large \color{green}{H = 10}\]

OpenStudy (anonymous):

it would be 10.